Scientific Notation in problem solving. Solve.
Coral Reefs. There are 10 million bacteria per square centimeter of coral in a coral reef. The coral reefs near the Hawaiian Islands cover 14,000 km^2. How many bacteria are there in Hawaii's coral reef?
My answer was 1.4 x 10^3. The books answer was 1.4 x 10^21. If I'm wrong, show me how it's done.

Answer by stanbon(75887)

Scientific Notation in problem solving. Solve.
Coral Reefs.
There are 10 million bacteria per square centimeter of coral in a coral reef.
10 million = 10*10^6* = 10^7 (bacteria)/cm^2
-----------------------------
The coral reefs near the Hawaiian Islands cover 14,000 km^2.
1 km^2 (100,000 cm/1 km)^2 = 10^10 cm^2
--------------------------
14000 km^2 = 1.4*10^3*(10^10)cm^2 = 1.4*(10^13)cm^2
----------------------------
How many bacteria are there in Hawaii's coral reef?
Ans: 1.4(10^13)*10*10^7 = 1.4*10^20 bacteria in the reef.
============
Cheers,
Stan H.
-----------